La bibliographie autour de la Méthode du Maximum d’Entropie

A. Coulais

4 avril 2003

Références

Abl74 [1] J. Ables, Maximum Entropy spectral analysis, A&A Supplement Series, 15:383–393, 1974, . Bev93 [2] R. Bevensee, Maximum Entropy solutions to scientific problems, PTR Prentice Hall, Englewood Cliffs, New Jersey

07632, 1993.

Bhan78 [3] R. Bhandari, Maximum Entropy spectral analysis- some comments, A&A, 70(0):331–333, mois inconnu 1978, . Buck91 [4] B. Buck and V. Macaulay, eds., Maximum Entropy in action, Oxford Science Publications, 1991. Fried72 [5] B. Frieden, Restoring with maximum likelihood and maximum entropy, Journal of the Optical Society of America,

62:511–518, 1972, .

Jaynes57 [6] E. Jaynes, Information theory and staitical mechanics, Phys. Rev., 106:620, 1957, .

KNN81 [7] M. Komesaroff, R. Narayan, and R. Nityananda, the Maximum Entropy Method of image restoration-

properties ans limitations, A&A, 93:269–281, 1981, .

LimMal81 [8] J. Lim and N. Malik, A new algorithm for two-dimensional maximum entropy power spectrum estimation, IEEE

Trans. on acoustics, speech, and signal processing, 29(3):401–413, Juin 1981, .

Rob84_NN [9] R. Narayan and R. Nityananda, Maximum Entropy : Flexilibity versus fundamentalism, in Indirect Imaging,

J. Roberts, ed., URSI, Australie 1983, 1984, pp. 281–290.

NN86 [10] R. Narayan and R. Nityananda, Maximum Entropy image restoration in astronomy, Annual Revue of A&A,

24:127–170, 1986, .

NN82 [11] R. Nityananda and R. Narayan, Maximum Entropy image reconstruction. a practical non-information theoretic

approach, J. of Astrophys. Astr., 3:419–450, 1982, .

Pon73 [12] Ponsonby, An Entropy measure of partially polarized radiation and its application to estimating radio sky polarization from incomplete aperture synthesis data by the maximum Entropy method, Monthly Notices of the Royal Astronomy Society, 163:369–380, 1973, .

Wer77 [13] S. Wernecke, Two-dimensional maximum entropy reconstruction of radio brightness, Radio science, 12(5):831–844, Septembre 1977, .